OpenStudy (anonymous):

For 0°≤θ≤360°, how many roots does the equation 3sin^(2)x+2sinx-1=0 have? let sinx=y => 3y^(2)+2y-1=0 y=1/3 or -1 .....

OpenStudy (raden):

3y^(2) + 2y - 1 = 0 this, can be factored (3y - 1)(y + 1) = 0 now solve for y

OpenStudy (anonymous):

y=1/3 or -1

OpenStudy (raden):

yup

OpenStudy (anonymous):

sinx=1/3 sinx=-1

OpenStudy (raden):

hmm.. for sin x = 1/3, this is not special angle in trigono. so use ur calculator to get x by take sin^-1

OpenStudy (anonymous):

After all, you are only asked how many roots, right? Or are they asking for specifics?

OpenStudy (anonymous):

@RadEn sinx=1/3 x=19.471220634491 @PeterPan i also want to know what're the roots

OpenStudy (anonymous):

Well, in that case, it doesn't look like there's much of a choice but to use calculators :)

OpenStudy (anonymous):

one of them is 19.47

OpenStudy (anonymous):

180-19.47=160.53 <<< is it correct?

OpenStudy (anonymous):

Way ahead of me :)

OpenStudy (anonymous):

any roots else?

OpenStudy (anonymous):

sin(x) = -1

OpenStudy (anonymous):

270

OpenStudy (anonymous):

Those are all, I suppose. Three roots.

OpenStudy (anonymous):

okay, thanks

OpenStudy (raden):

agreed ^

OpenStudy (anonymous):

and 0°≤θ≤360°, how many roots does the equation cos^(3)x=cosx have?

OpenStudy (anonymous):

Do the same, but first, rearrange into cos^3(x) - cos(x) = 0 and let y = cos(x) this time

OpenStudy (anonymous):

y^(3)-y=0

OpenStudy (anonymous):

Yes... and this can be factored. quite nicely, actually.

OpenStudy (anonymous):

It factors neatly into y(y - 1)(y + 1)

OpenStudy (anonymous):

So what are the values for y?

OpenStudy (anonymous):

cosx=0 cosx=1 cosx=-1

OpenStudy (anonymous):

That's right. And these are all special, aren't they? ^.^

OpenStudy (anonymous):

90 , 270, 0, 360, 180

OpenStudy (anonymous):

five??

OpenStudy (anonymous):

five roots... I find it a bit weird that both 0 and 360 are included, aren't they essentially the same? But yeah, you said your interval was 0°≤θ≤360° So you're correct. And yes, five.

OpenStudy (anonymous):

another person also confusing that 0 and 360, i think these two should be included

OpenStudy (anonymous):

It should be safe enough. ^.^

OpenStudy (anonymous):

For 0°≤θ≤360°, how many roots does the equation 2cos^(2)θ-5sinθ-4=0 have ?

OpenStudy (anonymous):

Leveling up? ^_^ Why did you change cos to sin? That's not allowed :)

OpenStudy (anonymous):

-2sin^(2)θ-5sinθ-2=0

OpenStudy (anonymous):

ohh. okay...

OpenStudy (anonymous):

2cos^(2)θ-5sinθ-4=0 2(1-sin^(2)θ)-5sinθ-4=0

OpenStudy (anonymous):

Then the answer is zero roots :D

OpenStudy (anonymous):

sinx=-2 or -0.5

OpenStudy (anonymous):

or that... I thought you said imaginary?

OpenStudy (anonymous):

nothing, i press the calculator wrongly

OpenStudy (anonymous):

Okay, so you've worked out that sinx = -2 or sin(x) = -0.5 What's your conclusion? :D

OpenStudy (anonymous):

i got 150° so far..

OpenStudy (anonymous):

Okay, just had to convince myself :D 150 is good, and there should be one more...

OpenStudy (anonymous):

wait, no.... 150 is not good at all, you need sin(x) = -0.5 sin(150) = 0.5

OpenStudy (anonymous):

210?

OpenStudy (anonymous):

Yeah, much better :) And one more.

OpenStudy (anonymous):

sin^(-1) -2?

OpenStudy (anonymous):

no... uhh one more value for x which would make sin(x) = -0.5 NOTE: There is no value for x which would make sin(x) = -2 sin only ranges from -1 to 1 ^.^

OpenStudy (anonymous):

ohh, yes..

OpenStudy (anonymous):

sinx=-0.5 x=-30 270-(-30)=300 ?

OpenStudy (anonymous):

Most excellent ^.^

OpenStudy (anonymous):

3 roots?

OpenStudy (anonymous):

Only two :)

OpenStudy (anonymous):

the two values for x which would make sin(x) = -0.5

OpenStudy (anonymous):

ohh yes.

OpenStudy (anonymous):

i understand the concept now :) thank you very much

OpenStudy (anonymous):

You're quite welcome ^.^

OpenStudy (anonymous):

how about...

OpenStudy (anonymous):

For 0°≤θ≤360°, how many roots does the equation tanx=2sinx have?